• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fansler

    Log entry time 15:50:26 on October 22, 2009

    Entry number 296120


    keyword=Cryo since power loss

    Cryo since power loss 10-20-09 All vacuums were checked, and verified good (0 Mtorr). We replaced the blower on the LQ1. The return valves (WEKA) on the left and right Q2/Q3 magnets were verified open and unplugged. The 04/16 valves on the Q1’s and Q2’s were calibrated. The calibration drifted due to radiation damage. Q3 valves were verified with in calibration. All lead flows were adjusted to “normal” levels. All magnet lead heaters were checked and verified operational. Lead flows may need to adjusted again after target cool down.